OpenStudy (anonymous):

what is the distance between (-6,-2)and(2,4) i picked 2square7

OpenStudy (amistre64):

if you can recall the pythag thrm; this is the same thing they give you a distance for 2 legs; and call them x and y find the hypotenuse or distance with the pythag thrm

OpenStudy (amistre64):

our first step should be to get the distance between xs and ys by subtracting one point from the other like this: (-6,-2) -( 2, 4) ------- -8,-6 ; so one leg = 8 and the other leg = 6 ; we ignore the - sign since direction is unimportant

OpenStudy (amistre64):

(8^2) + (6^2) = distance^2 64 + 36 = d^2 100 = d^2 ; what would our answer be for d?

OpenStudy (anonymous):

10

OpenStudy (anonymous):

10

OpenStudy (amistre64):

10 is correct :) good job

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

ok i got another one we can work out